OpenStudy (anonymous):

Find all solutions of the equation tan(x/2)+cos(x)=1

OpenStudy (anonymous):

Try the tangent half-angle formula.

OpenStudy (anonymous):

\[\Huge \tan \left(\frac{\theta}{2} \right)= \frac{\sin\theta}{1 + \cos\theta} = \frac{1-\cos\theta}{\sin\theta}\]

OpenStudy (anonymous):

I suggest trying out the second version.

OpenStudy (anonymous):

how would my equation fit into that?

OpenStudy (anonymous):

You have a \(\tan \frac{x}{2}\), so expand it through the identity.

jimthompson5910 (jim_thompson5910):

are you given exact or approximate solutions as answer choices?

OpenStudy (anonymous):

my options are: x=0 x=pi/2, pi x=0, pi/2 x=0, pi/2, pi x=0, pi

OpenStudy (anonymous):

If you're given options, I'd suggest just plugging in the values and seeing which ones work and which don't.

jimthompson5910 (jim_thompson5910):

Alright, thanks for giving the answer choices. It helped me see where I made a typo and I was able to fix it. tan(x/2) + cos(x) = 1 (sin(x))/(1+cos(x)) + cos(x) = 1 sin(x) + cos(x)(1+cos(x)) = 1+cos(x) sin(x) + cos(x) + cos^2(x) = 1+cos(x) sin(x) + cos(x) + cos^2(x) - 1 - cos(x) = 0 sin(x) + cos^2(x) - 1 = 0 sin(x) + 1 - sin^2(x) - 1 = 0 sin(x) - sin^2(x) = 0 -sin^2(x) + sin(x) = 0 -1(sin^2(x) - sin(x)) = 0 sin^2(x) - sin(x) = 0 sin(x)(sin(x) - 1) = 0 sin(x) = 0 or sin(x)-1 = 0 sin(x) = 0 or sin(x) = 1 Solve sin(x) = 0: sin(x) = 0 x = arcsin(0) x = 0 ----------------- Now solve sin(x) = 1 sin(x) = 1 x = arcsin(1) x = pi/2 So the solutions are x=0, pi/2